OpenStudy (anonymous):

8^8*8^4 What is each expression written using each base only once?

OpenStudy (anonymous):

8^32

OpenStudy (anonymous):

recall: a^n*a^m=a^(n*m)

OpenStudy (anonymous):

lagrange

OpenStudy (anonymous):

No. It is a^(n+m). So, 8^12.

OpenStudy (anonymous):

thx lagrange

OpenStudy (anonymous):

(a^n)^m = a^(nm)

OpenStudy (anonymous):

i used that answer and got it wrong? choices are A:512^12 B:8^32 C:24^12 D:8^12

OpenStudy (anonymous):

In this case D is the right answer.

OpenStudy (anonymous):

What is the simplified form of the expression?

OpenStudy (anonymous):

Post separately.

OpenStudy (anonymous):